# Matrix product state approach to lossy boson sampling and noisy IQP sampling

### v1

@@ I Introduction
It is widely believed that a quantum computer can efficiently solve certain problems that are hard for classical computers, such as integer factoring and Hamiltonian simulation [1, 2]. Since one of the most critical obstacles to realizing a quantum computer is noise, quantum error correction (QEC) has been developed to actively detect and correct errors, enabling fault-tolerant quantum computation [3, 4]. Despite substantial advances in both quantum hardware and the theory and implementation of QEC [5, 6], currently available quantum devices cannot yet fully exploit the QEC technique in a fully scalable manner. Consequently, extensive efforts have focused on achieving quantum advantage using currently available quantum devices, often referred to as noisy intermediate-scale quantum (NISQ) devices [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17].
@@
To pursue quantum advantage with NISQ devices, quantum sampling problems have emerged as promising candidates due to their complexity-theoretic hardness results in noiseless cases and their relatively feasible experimental requirements, compared with more sophisticated quantum algorithms [7, 8, 9]. The most representative examples of sampling problems include boson sampling, random circuit sampling, and instantaneous quantum polynomial-time (IQP) sampling [7, 18, 19]. In fact, the experimental implementation of these sampling problems has been conducted to claim quantum advantage by experiments [20, 12, 14, 9, 21, 17].
@@
However, such claims have been continuously challenged due to the noise in the realistic experiment (e.g., Refs. [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]). The main insight for this is that physical noise often degrades and ultimately destroys quantum advantage. For example, when a depolarizing noise occurs in quantum circuits and the circuit depth is large, the system accumulates entropy, causing the output state to converge to the maximally mixed state, which is easily simulated classically [36]. Similarly, in photonic systems, it has been shown that photon loss makes a quantum state converge to a thermal state, which is easy to classically simulate [25, 26]. As such, physical noise may make sampling problems that are hard to classically simulate when noiseless become easy to simulate. Hence, to rigorously assess the potential advantage of quantum devices, it is essential to understand whether and when sampling problems become easy to classically simulate under realistic noise conditions.
@@
Focusing specifically on boson sampling, one of the most widely studied sampling models, as aforementioned, when the loss rate is sufficiently large, boson sampling becomes classically easy. Motivated by this observation, a variety of classical simulation algorithms have been developed [37, 25, 38, 39, 27, 32]. One line of work [25] used a strategy to approximate a lossy quantum state by a quantum state that is classically easy to simulate. For example, Ref. [25] and Ref. [24] showed that a lossy single photon state can be approximated by a thermal state or a separable state, respectively, when the loss rate is high. Furthermore, this idea has been extended to lossy Gaussian boson sampling (GBS) [26]. Nonetheless, when applied to recent GBS experiments, while it provides an important benchmark, the classical algorithm was not sufficient to simulate them [14, 12]. A key limitation of these approaches is that, since they rely on approximating a given state by the nearest classically simulable state, their performance cannot be systematically improved by allocating additional computational resources. Hence, below a loss rate threshold, the algorithm fails to operate.
@@
Notably, a more recently proposed classical algorithm based on the matrix product state (MPS) has overcome this limitation and successfully simulated the state-of-the-art GBS experiments at the moment. The crucial feature of the new classical algorithm is the tunability of its performance by providing more computational resources, which translates to the bond dimension of the MPS. However, because the method in Ref. [33] relies on the Gaussianity of the quantum system, it has not been extended to other systems so far, despite its high performance in practice.
@@
In this work, we generalize the MPS for GBS proposed in Ref. [33] to simulate lossy boson sampling and noisy IQP sampling. In other words, we develop MPS-based classical algorithms for lossy boson sampling and noisy IQP sampling that correspond to the state-of-the-art classical algorithm for lossy GBS; hence, our classical algorithms also enjoy the property that the accuracy is tunable by consuming computational resources, unlike the previous algorithms [25, 24]. The central idea of our approach is to construct a pure-state decomposition of a lossy or noisy input state such that each component remains weakly entangled after circuit evolution. For the boson sampling case, we consider various input states, such as single-photon states, Fock states, and cat states as input and analyze the performance of our algorithm using the entanglement entropy [40, 41]. Notably, we find a similar classically simulable range of the transmission rate to GBS and previous results [33, 24, 25], which is given as $O(1/\sqrt{N})$, where $N$ is the input photon number. We then estimate the bond dimension required to simulate, providing guidelines for future experiments. We further extend our framework to noisy IQP sampling and then demonstrate that the performance of our classical algorithm is comparable to that of the recently proposed algorithm [42].
@@ II.1 Boson sampling
Boson sampling is a sampling problem that is proposed to demonstrate quantum advantage using photons [7]. To formally introduce boson sampling, consider an $M$-mode bosonic system and let $\hat{a}_{i}$ and $\hat{a}_{i}^{\dagger}$ denote the bosonic annihilation and creation operators of the $i$th mode, respectively, satisfying the canonical commutation relations, $\hat{a}_{i},\hat{a}_{j}^{\dagger}]=\delta_{ij}$ and $\hat{a}_{i},\hat{a}_{j}]=0$, where $1\leq i\leq M$. The input state of boson sampling is $N$ single photons in the first $N$ modes out of a total of $M$ modes, with the rest of the modes being initialized as vacuum, i.e., the input state is written as
@@
where $\hat{b}_{i}^{\dagger}$ represents the creation operator of the $i$th output mode. Here, $U$ is often chosen to follow the Haar-random unitary ensemble for hardness results [7]. Such a Haar-random linear optical circuit can be implemented using two-mode beam splitters and phase shifters, with the circuit depth being linear in $M$ [43]. After the circuit, the output state is measured by photon number-resolving detectors, which give an output sample $(t_{1},\dots,t_{M})\in\mathbb{Z}_{\geq 0}^{M}$, where $t_{i}$ represents the output photon number on the $i$th output mode, satisfying $\sum_{i=1}^{M}t_{i}=N$ since linear-optical circuits preserve the photon number.
Boson sampling has attracted considerable attention due to its hardness result. More specifically, lossless boson sampling is proven to be hard to classically simulate under some plausible conjectures [7]. The main observation for the hardness result is that the corresponding output probability is described by a matrix function, so-called the permanent:
@@
Although the ideal boson sampling is believed to be hard to simulate classically, the actual experiments that implement boson sampling suffer from practical imperfections that hinder the demonstration of quantum advantage. In particular, one of the dominant imperfections is photon loss in the current experiments [44], and many studies demonstrate that under photon loss, photon sampling becomes easier to classically simulate [25, 24, 27, 26, 32, 30]. Therefore, it is crucial to consider the effect of photon loss and understand its significance for rigorously evaluating the classical simulability of experimental boson sampling.
@@
Throughout this work, we assume that the optical devices have the same transmission rate and each mode has the same depth for simplicity; therefore, the transmission rate and the loss rate of each mode are assumed to be uniform. This uniform loss assumption allows each layer of the loss channel to commute with all other linear optical devices and be brought to the forefront of the lossless circuit. Consequently, the model reduces to a Haar random interferometer preceded by a uniform loss channel, as shown in Fig. 1. In the case of nonuniform loss, we can generalize this method by commuting the uniform loss part to the front and setting the overall transmission rate as the maximum among the modes in the remaining optics, and commuting to the front (see Ref. [24] for more details).
@@ II.2 IQP sampling
IQP sampling is another sampling problem based on qubits, which is also widely believed to be computationally intractable for classical computers [19]. The $n$-qubit IQP sampling operates as follows. First, the input state is prepared as $|0\rangle^{\otimes n}$ and then evolves under a quantum circuit. At first, a layer of $n$ Hadamard gates is applied to the initial state $|0\rangle^{\otimes n}$, which transforms the state to the uniform superposition of all bitstrings with length $n$:
@@
where $f$ is a real function determined by the diagonal gates. Thus, the diagonal gates change the relative phases between different computational basis elements. Finally, a layer of Hadamard gates is applied after diagonal gates, and the output state is measured in the $Z$ basis. Like boson sampling, IQP sampling is believed to be classically intractable: if it admitted an efficient classical simulation, the polynomial hierarchy would collapse to the third level, a consequence widely considered implausible [19].
As for the boson sampling case, however, many recent results show that IQP sampling may also be efficiently simulated when noise occurs in the system [45, 42, 46]. In this work, we exploit our MPS method to investigate the effect of noise on the hardness of IQP circuits and compare with other existing methods. To investigate the effect of noise on the complexity of simulating IQP sampling, we consider Pauli noises in IQP circuits, which is a dominant noise model in many physical devices. A Pauli noise channel transforms a quantum state $\hat{\rho}$ as follows:
@@ III MPS and classical simulability
As mentioned in the introduction, this work focuses on the MPS method for classical simulation. Thus, we provide a brief review of the MPS, which is particularly useful when a given quantum state is slightly entangled [41]. More specifically, consider a quantum state $|\psi\rangle$ written as
@@
In particular, $|\psi\rangle$ is classically approximable by MPS, i.e., the required bond dimension is at most polynomial in $L$, when $S_{\alpha}^{l}(|\psi\rangle)=O(\log L)$ for all $l$-site reduced states of $|\psi\rangle$ [41]. Using this as a sufficient simulability condition and relating it to the photon loss rate or Pauli noise rate, we can obtain the simulable range of loss or noise, respectively.
On the other hand, it is also well-known that if the ERE scales algebraically, i.e., $S_{\alpha}^{l}(|\psi\rangle)=\Omega(L^{\kappa})$ and $\kappa>0$ for some $l$, the corresponding state cannot be efficiently approximated by an MPS [41]. Using this, we also characterize the regime of loss or noise that is hard to simulate using MPS.
@@ IV.1 MPS for simulating noisy quantum systems
As emphasized in Sec. III, the MPS is particularly useful for simulating low-entangled quantum systems. Therefore, the MPS method can be treated as a suitable choice to simulate noisy quantum systems because noise typically destroys the entanglement. However, an immediate obstacle to applying the MPS method for simulating noisy quantum systems, such as lossy boson sampling, is that the noisy quantum states are not pure in general, while the MPS method supports only pure states. Thus, the MPS method’s restriction to pure states hinders us from exploiting its feature for simulating noisy systems. An interesting way to overcome this limitation is to vectorize the density matrix so that we can describe a mixed state by a pure state formalism, which is called the MPO method [47, 27, 32]. Indeed, this approach exhibits that as the system becomes noisier, it is easier to perform classical simulation. However, compared to the MPS method, the MPO method requires more bond dimensions and more resources for updating the state, resulting in a greater resource requirement.
More recently, another method has been proposed to overcome this limitation, simulating state-of-the-art GBS experiments classically [33]. The main idea of this method is to decompose a noisy mixed state into a probabilistic mixture of pure states, each of which has low entanglement as the system becomes noisier. Such a method is the state-of-the-art classical simulator for GBS, to the best of our knowledge. Therefore, its generalization to other noisy quantum systems may promise to find one of the best classical algorithms at this moment. Nevertheless, the proposed decomposition appears to be highly dependent on the Gaussianity of the state, and thus, its generalization to more general cases has not been conducted. We now generalize this method to lossy boson sampling and noisy IQP sampling by finding an appropriate decomposition of noisy quantum states, each of which has low entanglement.
@@ IV.2 Decomposition of lossy input state
Instead of employing the MPO method, which requires substantial computational resources, we adopt the MPS formalism to simulate mixed states. In this approach, a mixed state is expressed as a convex combination of pure states, and each pure state is then simulated according to its corresponding probability weight [33]. More specifically, if we have a mixed state input $\hat{\rho}_{\text{in}}=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ and the rest of the quantum circuit is unitary $\hat{U}$, one may simply sample $i$ from the probability distribution $\{p_{i}\}$ and then simulate $\hat{U}|\psi_{i}\rangle$. Such a process obviously exactly simulates the sampling from the density matrix $\hat{\rho}=\sum_{i}p_{i}\hat{U}|\psi_{i}\rangle\langle\psi_{i}|\hat{U}^{\dagger}$. A subtlety with this approach is that, in principle, many different decompositions are consistent with the given density matrix, and different decompositions may incur different computational costs. An obvious optimal choice for an efficient MPS simulation is a decomposition such that each of the pure-state components $|\psi_{i}\rangle$ has a small entanglement throughout the circuit.
For this particular setup of lossy boson sampling, one naive but natural choice of the decomposition of $\hat{\sigma}$ is a convex sum of $|0\rangle\langle 0|$ and $|1\rangle\langle 1|$ with probabilities $1-\eta$ and $\eta$, respectively. However, one may easily notice a drawback of this decomposition, which leads to a significant computational cost [31]. More specifically, since the transmission rate is $\eta$, the sampled pure state input contains an average $\eta N$ number of input $|1\rangle\langle 1|$. To simulate this sampling problem in polynomial time, $\eta N=O(\log N)$ is required because the complexity of boson sampling is exponential to the number of single photon inputs. Therefore, the simulable range of $\eta$ is too limited compared to the known simulable range $\eta=O(1/\sqrt{N})$ [25].

### v2

@@ I Introduction
It is widely believed that a quantum computer can efficiently solve certain problems that are hard for classical computers, such as integer factoring and Hamiltonian simulation [shor1994algorithms, lloyd1996universal]. Since one of the most critical obstacles to realizing a quantum computer is noise, quantum error correction (QEC) has been developed to actively detect and correct errors, enabling fault-tolerant quantum computation [shor1995scheme, steane1996error]. Despite substantial advances in both quantum hardware and the theory and implementation of QEC [google2025quantum, terhal2015quantum], currently available quantum devices cannot yet fully exploit the QEC technique in a fully scalable manner. Consequently, extensive efforts have focused on achieving quantum advantage using currently available quantum devices, often referred to as noisy intermediate-scale quantum (NISQ) devices [aaronson2011computational, bouland2019complexity, arute2019quantum, wu2021strong, morvan2023phase, zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian, young2024atomic, liu2025robust].
@@
To pursue quantum advantage with NISQ devices, quantum sampling problems have emerged as promising candidates due to their complexity-theoretic hardness results in noiseless cases and their relatively feasible experimental requirements, compared with more sophisticated quantum algorithms [aaronson2011computational, bouland2019complexity, arute2019quantum]. The most representative examples of sampling problems include boson sampling, random circuit sampling, and instantaneous quantum polynomial-time (IQP) sampling [aaronson2011computational, boixo2018characterizing, bremner2011classical]. In fact, the experimental implementation of these sampling problems has been conducted to claim quantum advantage by experiments [wang2019boson, zhong2020quantum, madsen2022quantum, arute2019quantum, bluvstein2024logical, liu2025robust].
@@
However, such claims have been continuously challenged due to the noise in the realistic experiment (e.g., Refs. [neville2017classical, clifford2018classical, oszmaniec2018classical, garcia2019simulating, qi2020regimes, oh2021classical, quesada2022quadratic, bulmer2022boundary, oh2022classical, oh2023classical, liu2023simulating, oh2024classical, oh2025recent, oh2025classical]). The main insight for this is that physical noise often degrades and ultimately destroys quantum advantage. For example, when a depolarizing noise occurs in quantum circuits and the circuit depth is large, the system accumulates entropy, causing the output state to converge to the maximally mixed state, which is easily simulated classically [aharonov1996limitations]. Similarly, in photonic systems, it has been shown that photon loss makes a quantum state converge to a thermal state, which is easy to classically simulate [garcia2019simulating, qi2020regimes]. As such, physical noise may make sampling problems that are hard to classically simulate when noiseless become easy to simulate. Hence, to rigorously assess the potential advantage of quantum devices, it is essential to understand whether and when sampling problems become easy to classically simulate under realistic noise conditions.
@@
Focusing specifically on boson sampling, one of the most widely studied sampling models, as aforementioned, when the loss rate is sufficiently large, boson sampling becomes classically easy. Motivated by this observation, a variety of classical simulation algorithms have been developed [rahimi2016sufficient, garcia2019simulating, huang2019simulating, clifford2020faster, oh2021classical, liu2023simulating]. One line of work [garcia2019simulating] used a strategy to approximate a lossy quantum state by a quantum state that is classically easy to simulate. For example, Ref. [garcia2019simulating] and Ref. [oszmaniec2018classical] showed that a lossy single photon state can be approximated by a thermal state or a separable state, respectively, when the loss rate is high. Furthermore, this idea has been extended to lossy Gaussian boson sampling (GBS) [qi2020regimes]. Nonetheless, when applied to recent GBS experiments, while it provides an important benchmark, the classical algorithm was not sufficient to simulate them [madsen2022quantum, zhong2020quantum]. A key limitation of these approaches is that, since they rely on approximating a given state by the nearest classically simulable state, their performance cannot be systematically improved by allocating additional computational resources. Hence, below a loss rate threshold, the algorithm fails to operate.
@@
Notably, a more recently proposed classical algorithm based on the matrix product state (MPS) has overcome this limitation and successfully simulated the state-of-the-art GBS experiments at the moment. The crucial feature of the new classical algorithm is the tunability of its performance by providing more computational resources, which translates to the bond dimension of the MPS. However, because the method in Ref. [oh2024classical] relies on the Gaussianity of the quantum system, it has not been extended to other systems so far, despite its high performance in practice.
@@
In this work, we generalize the MPS for GBS proposed in Ref. [oh2024classical] to simulate lossy boson sampling and noisy IQP sampling. In other words, we develop MPS-based classical algorithms for lossy boson sampling and noisy IQP sampling that correspond to the state-of-the-art classical algorithm for lossy GBS; hence, our classical algorithms also enjoy the property that the accuracy is tunable by consuming computational resources, unlike the previous algorithms [garcia2019simulating, oszmaniec2018classical]. The central idea of our approach is to construct a pure-state decomposition of a lossy or noisy input state such that each component remains weakly entangled after circuit evolution. For the boson sampling case, we consider various input states, such as single-photon states, Fock states, and cat states as input and analyze the performance of our algorithm using the entanglement entropy [vidal2003efficient, schuch2008entropy]. Notably, we find a similar classically simulable range of the transmission rate to GBS and previous results [oh2024classical, oszmaniec2018classical, garcia2019simulating], which is given as $O(1/\sqrt{N})$, where $N$ is the input photon number. We then estimate the bond dimension required to simulate, providing guidelines for future experiments. We further extend our framework to noisy IQP sampling and then demonstrate that the performance of our classical algorithm is comparable to that of the recently proposed algorithm [nelson2024polynomial].
@@ II.1 Boson sampling
Boson sampling is a sampling problem that is proposed to demonstrate quantum advantage using photons [aaronson2011computational]. To formally introduce boson sampling, consider an $M$-mode bosonic system and let $\hat{a}_{i}$ and $\hat{a}_{i}^{\dagger}$ denote the bosonic annihilation and creation operators of the $i$th mode, respectively, satisfying the canonical commutation relations, $\hat{a}_{i},\hat{a}_{j}^{\dagger}]=\delta_{ij}$ and $\hat{a}_{i},\hat{a}_{j}]=0$, where $1\leq i\leq M$. The input state of boson sampling is $N$ single photons in the first $N$ modes out of a total of $M$ modes, with the rest of the modes being initialized as vacuum, i.e., the input state is written as
@@
where $\hat{b}_{i}^{\dagger}$ represents the creation operator of the $i$th output mode. Here, $U$ is often chosen to follow the Haar-random unitary ensemble for hardness results [aaronson2011computational]. Such a Haar-random linear optical circuit can be implemented using two-mode beam splitters and phase shifters, with the circuit depth being linear in $M$ [russell2017direct]. After the circuit, the output state is measured by photon number-resolving detectors, which give an output sample $(t_{1},\dots,t_{M})\in\mathbb{Z}_{\geq 0}^{M}$, where $t_{i}$ represents the output photon number on the $i$th output mode, satisfying $\sum_{i=1}^{M}t_{i}=N$ since linear-optical circuits preserve the photon number.
Boson sampling has attracted considerable attention due to its hardness result. More specifically, lossless boson sampling is proven to be hard to classically simulate under some plausible conjectures [aaronson2011computational]. The main observation for the hardness result is that the corresponding output probability is described by a matrix function, so-called the permanent:
@@
Although the ideal boson sampling is believed to be hard to simulate classically, the actual experiments that implement boson sampling suffer from practical imperfections that hinder the demonstration of quantum advantage. In particular, one of the dominant imperfections is photon loss in the current experiments [wang2018toward], and many studies demonstrate that under photon loss, photon sampling becomes easier to classically simulate [garcia2019simulating, oszmaniec2018classical, oh2021classical, qi2020regimes, liu2023simulating, oh2022classical]. Therefore, it is crucial to consider the effect of photon loss and understand its significance for rigorously evaluating the classical simulability of experimental boson sampling.
@@
Throughout this work, we assume that the optical devices have the same transmission rate and each mode has the same depth for simplicity; therefore, the transmission rate and the loss rate of each mode are assumed to be uniform. This uniform loss assumption allows each layer of the loss channel to commute with all other linear optical devices and be brought to the forefront of the lossless circuit. Consequently, the model reduces to a Haar random interferometer preceded by a uniform loss channel, as shown in Fig. 1. In the case of nonuniform loss, we can generalize this method by commuting the uniform loss part to the front and setting the overall transmission rate as the maximum among the modes in the remaining optics, and commuting to the front (see Ref. [oszmaniec2018classical] for more details).
@@ II.2 IQP sampling
IQP sampling is another sampling problem based on qubits, which is also widely believed to be computationally intractable for classical computers [bremner2011classical]. The $n$-qubit IQP sampling operates as follows. First, the input state is prepared as $|0\rangle^{\otimes n}$ and then evolves under a quantum circuit. At first, a layer of $n$ Hadamard gates is applied to the initial state $|0\rangle^{\otimes n}$, which transforms the state to the uniform superposition of all bitstrings with length $n$:
@@
where $f$ is a real function determined by the diagonal gates. Thus, the diagonal gates change the relative phases between different computational basis elements. Finally, a layer of Hadamard gates is applied after diagonal gates, and the output state is measured in the $Z$ basis. Like boson sampling, IQP sampling is believed to be classically intractable: if it admitted an efficient classical simulation, the polynomial hierarchy would collapse to the third level, a consequence widely considered implausible [bremner2011classical].
As for the boson sampling case, however, many recent results show that IQP sampling may also be efficiently simulated when noise occurs in the system [bremner2017achieving, nelson2024polynomial, rajakumar2025polynomial]. In this work, we exploit our MPS method to investigate the effect of noise on the hardness of IQP circuits and compare with other existing methods. To investigate the effect of noise on the complexity of simulating IQP sampling, we consider Pauli noises in IQP circuits, which is a dominant noise model in many physical devices. A Pauli noise channel transforms a quantum state $\hat{\rho}$ as follows:
@@ III MPS and classical simulability
As mentioned in the introduction, this work focuses on the MPS method for classical simulation. Thus, we provide a brief review of the MPS, which is particularly useful when a given quantum state is slightly entangled [schuch2008entropy]. More specifically, consider a quantum state $|\psi\rangle$ written as
@@
In particular, $|\psi\rangle$ is classically approximable by MPS, i.e., the required bond dimension is at most polynomial in $L$, when $S_{\alpha}^{l}(|\psi\rangle)=O(\log L)$ for all $l$-site reduced states of $|\psi\rangle$ [schuch2008entropy]. Using this as a sufficient simulability condition and relating it to the photon loss rate or Pauli noise rate, we can obtain the simulable range of loss or noise, respectively.
On the other hand, it is also well-known that if the ERE scales algebraically, i.e., $S_{\alpha}^{l}(|\psi\rangle)=\Omega(L^{\kappa})$ and $\kappa>0$ for some $l$, the corresponding state cannot be efficiently approximated by an MPS [schuch2008entropy]. Using this, we also characterize the regime of loss or noise that is hard to simulate using MPS.
@@ IV.1 MPS for simulating noisy quantum systems
As emphasized in Sec. III, the MPS is particularly useful for simulating low-entangled quantum systems. Therefore, the MPS method can be treated as a suitable choice to simulate noisy quantum systems because noise typically destroys the entanglement. However, an immediate obstacle to applying the MPS method for simulating noisy quantum systems, such as lossy boson sampling, is that the noisy quantum states are not pure in general, while the MPS method supports only pure states. Thus, the MPS method’s restriction to pure states hinders us from exploiting its feature for simulating noisy systems. An interesting way to overcome this limitation is to vectorize the density matrix so that we can describe a mixed state by a pure state formalism, which is called the MPO method [noh2020efficient, oh2021classical, liu2023simulating]. Indeed, this approach exhibits that as the system becomes noisier, it is easier to perform classical simulation. However, compared to the MPS method, the MPO method requires more bond dimensions and more resources for updating the state, resulting in a greater resource requirement.
More recently, another method has been proposed to overcome this limitation, simulating state-of-the-art GBS experiments classically [oh2024classical]. The main idea of this method is to decompose a noisy mixed state into a probabilistic mixture of pure states, each of which has low entanglement as the system becomes noisier. Such a method is the state-of-the-art classical simulator for GBS, to the best of our knowledge. Therefore, its generalization to other noisy quantum systems may promise to find one of the best classical algorithms at this moment. Nevertheless, the proposed decomposition appears to be highly dependent on the Gaussianity of the state, and thus, its generalization to more general cases has not been conducted. We now generalize this method to lossy boson sampling and noisy IQP sampling by finding an appropriate decomposition of noisy quantum states, each of which has low entanglement.
@@ IV.2 Decomposition of lossy input state
Instead of employing the MPO method, which requires substantial computational resources, we adopt the MPS formalism to simulate mixed states. In this approach, a mixed state is expressed as a convex combination of pure states, and each pure state is then simulated according to its corresponding probability weight [oh2024classical]. More specifically, if we have a mixed state input $\hat{\rho}_{\text{in}}=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ and the rest of the quantum circuit is unitary $\hat{U}$, one may simply sample $i$ from the probability distribution $\{p_{i}\}$ and then simulate $\hat{U}|\psi_{i}\rangle$. Such a process obviously exactly simulates the sampling from the density matrix $\hat{\rho}=\sum_{i}p_{i}\hat{U}|\psi_{i}\rangle\langle\psi_{i}|\hat{U}^{\dagger}$. A subtlety with this approach is that, in principle, many different decompositions are consistent with the given density matrix, and different decompositions may incur different computational costs. An obvious optimal choice for an efficient MPS simulation is a decomposition such that each of the pure-state components $|\psi_{i}\rangle$ has a small entanglement throughout the circuit.
For this particular setup of lossy boson sampling, one naive but natural choice of the decomposition of $\hat{\sigma}$ is a convex sum of $|0\rangle\langle 0|$ and $|1\rangle\langle 1|$ with probabilities $1-\eta$ and $\eta$, respectively. However, one may easily notice a drawback of this decomposition, which leads to a significant computational cost [oh2023classical]. More specifically, since the transmission rate is $\eta$, the sampled pure state input contains an average $\eta N$ number of input $|1\rangle\langle 1|$. To simulate this sampling problem in polynomial time, $\eta N=O(\log N)$ is required because the complexity of boson sampling is exponential to the number of single photon inputs. Therefore, the simulable range of $\eta$ is too limited compared to the known simulable range $\eta=O(1/\sqrt{N})$ [garcia2019simulating].